To prove that 2ⁿ>1+n√(2ⁿ⁻¹)∀n>2, using:
Cauchy-Schwarz Inequality;
Weierstrass Inequalities;
Tchebychev's Inequalities.
1
Expert's answer
2019-07-11T09:04:41-0400

Proof by Chebyshev's sum inequality.

This inequality says, that if a1<=a2<= ... <=an and b1>=b2>= ... >=bn, then

n * (a1 * b1 + ... + an * bn) <= (a1 + ... + an) * (b1 + ... + bn).

Let bi = 2n-i and ai = 2i-n for each i from [1, n].

It is easy to see, that a is increasing sequence and b is decreasing sequence.

Also ai * bi = 1 for each i from [1, n].

So after using Chebyshev's sum inequality:

n * (1 + ... + 1) <= (21-n + ... + 1) * (2n-1 + ... + 1), or

n2 <= (21-n + ... + 1) * (2n-1 + ... + 1).

Let's multiply each side of inequality by 2n-1:

n2 * 2n-1 <= (1 + ... + 2n-1) * (2n-1 + ... + 1), or

(n * 2(n-1) / 2)2 <= (1 + ... + 2n-1)2, or

n * 2(n-1) / 2 <= 1 + ... + 2n-1, because each side is greater than 0.

It is famous, that 20 + 21 + ... + 2n-1 = 2n - 1 (use Binomial Theorem, for example).

So:

n * 2(n-1) / 2 <= 2n - 1, or

n * 2(n-1) / 2 + 1 <= 2n.

It is important to say, that equality is impossible, because if n is even number, so left side isn't natural as right side, and if n is odd, then the right side is even and left side is odd.

So:

n * 2(n-1) / 2 + 1 < 2n.

Proof using Cauchy-Schwarz inequality.

This inequality says, that for sequences a and b inequaltity

(a1+ ... + an2) * (b12 + ... + bn2) >= (a1 * b1 + ... + an * bn)2 is always correct.

Let ai = 2(1 - i) / 2 and bi = 2(i - 1) / 2 for each i from [1, n].

Also ai * bi = 1 for each i from [1, n].

So after using Cauchy-Schwarz inequality:

(1 + 2-1 + ... + 2-n+1) * (1 + 21 + ... + 2n-1) >= n2, or

(2n - 1)2 / 2n-1 >=n2, or

2n - 1 >= n * 2(n -1) / 2.

But this inequality was already proved.

Proof using Weierstrass Inequalities.

This inequality says that for sequence an inequality

(1 + a1) * ... * (1 + an) >= 1 + a1 + ... + an is correct, where ai is [0; 1].

Let's rewrite inequality, which is needed to prove:

2n > 1 + n * 2(n - 1) / 2, or

(2(n-1) / 2)2 * 2 > 1 + n * 2(n - 1) / 2, or

2(n - 1) / 2 * (2(n + 1) / 2 - n) > 1.

Else it is enough to show, that 2(n + 1) / 2 >= n + 1.

Put in Weierstrass Inequalities a = {1, 1, ..., 1}.

2>= 1 + n, or 2n-1 >= n and 2n - i >= n + 1 - i (just decrease n i times).

Let t = (n + 1) / 2.

If n is odd, then 2t - 1 >= t, and 2t >= 2*t, or 2(n + 1) / 2 >= n + 1.

If n is even, then we can prove this inequality for n - 1 and using, that 2(n + 1) / 2 > 2n / 2, we could exclaim, that 2(n + 1) / 2 >= n + 1 (because step is equal 1).